John dissolves .5g of a white powder in 25g of benzene (FP 5oC) (kf benzene is 5.1) and finds the solution freezes at 3.7oC. Determine the MW of the compound!
1
Expert's answer
2020-04-06T12:05:40-0400

let the molar weight of the compound be "M".

now, "1.3=5.1\times\frac{\frac{5}{M}}{25\times10^{-3}}"

"M=784.615\frac{g}{mol}"